# GOT-Edit: Geometry-Aware Generic Object Tracking via Online Model Editing

### v1

@@ Abstract
Human perception for effective object tracking in a 2D video stream arises from the implicit use of prior 3D knowledge combined with semantic reasoning. In contrast, most generic object tracking (GOT) methods primarily rely on 2D features of the target and its surroundings while neglecting 3D geometric cues, which makes them susceptible to partial occlusion, distractors, and variations in geometry and appearance. To address this limitation, we introduce GOT-Edit, an online cross-modality model editing approach that integrates geometry-aware cues into a generic object tracker from a 2D video stream. Our approach leverages features from a pre-trained Visual Geometry Grounded Transformer to enable geometric cue inference from only a few 2D images. To tackle the challenge of seamlessly combining geometry and semantics, GOT-Edit performs online model editing with null-space constrained updates that incorporate geometric information while preserving semantic discrimination, yielding consistently better performance across diverse scenarios. Extensive experiments on multiple GOT benchmarks demonstrate that GOT-Edit achieves superior robustness and accuracy, particularly under occlusion and clutter, establishing a new paradigm for combining 2D semantics with 3D geometric reasoning for generic object tracking. Code is available [^1] https://github.com/chenshihfang/GOT.
@@ 1 Introduction
While geometric information is potentially beneficial for GOT, effectively balancing its contribution with crucial or even dominant semantic information remains a key challenge. As evidenced by our later experiment, a naive fusion strategy improves geometry attributes in tracking but degrades semantic attributes. To address this issue, we propose a novel online model editing technique that better integrates 3D geometric features from VGGT with 2D semantic features (Oquab et al., 2023) for GOT. Our approach is inspired by the null-space model editing from AlphaEdit (Fang et al., 2025), which is designed to introduce new knowledge into the null space of a trained model while preserving the semantic knowledge for optimal performance. However, AlphaEdit performs offline model editing, whereas GOT requires online updates to handle dynamically varying targets and backgrounds in both seen and unseen scenarios. To bridge this gap, we develop an online editing technique that enables a tracker to adaptively complement 2D semantics with 3D geometric features.
As illustrated in Figure 1, our system begins by extracting both semantic and geometric features from the current and reference frames. These features are then aligned and fused to create an enriched representation, which serves as new knowledge for online tracker adaptation. Built upon the ToMP (Mayer et al., 2022), our approach employs two model predictors: one for the semantic branch and one for the geometric branch. These predictors generate the model weights for the localization head. During tracking, the reference labels, which provide correspondences for the reference frames and act as few-shot examples of previously predicted and observed information, are dynamically updated to guide a tracker toward the target object. This process guides the model predictors to forecast model weights for the current frame in an online manner. Namely, the semantic model predictor estimates the semantic weights, while the geometry predictor generates complementary weights. A null-space constraint is applied before combining these two sets of weights to preserve the semantic information. Finally, the combined model weights are used by the localization head to localize the target in the current frame.
@@
Figure 1: The GOT-Edit Framework. GOT-Edit facilitates the understanding of 3D geometry to aid generic object tracking from 2D streaming inputs. It predicts semantic and geometric weights concurrently to incrementally adapt the tracking model. Through online model editing, it ensures geometry-aware, semantic-preserving updates to the tracking model. The solid red box marks the ground-truth target in the input reference frames. The dashed red boxes indicate these same annotations utilized for the online knowledge update within the geometry branch. The green box represents the final predicted tracking result.
@@ 3.2 GOT-Edit
This stabilized projector is then utilized in Eq. 10 to compute the geometry-aware perturbation weights.
@@ 4.1 Experimental Setting
Implementation Details. Our method is implemented using PyTorch 2.0.0 and CUDA 11.7. We train the model on eight NVIDIA RTX 4090 GPUs (24 GB each). DeepSpeed (Rasley et al., 2020) is integrated to accelerate training. We also verify that applying activation checkpointing to the tracker further reduces memory consumption, enabling training of the tracker at high resolution (378 $\times$ 378) on four 24 GB GPUs.
Inference is performed on a single NVIDIA RTX 4090 GPU and consumes approximately 9 GB of GPU memory during evaluation.
Following PiVOT (Chen et al., 2025a) and LoRAT (Lin et al., 2024), we utilize ViT-L as the backbone for image feature extraction, using weights pretrained with DINOv2 (Oquab et al., 2023). The backbone remains frozen during training with the tracker. For integrating geometric information, we extract intermediate features from the DPT head of VGGT (Wang et al., 2025a), which is kept frozen during training. Similar to PiVOT (Chen et al., 2025a), the model predictors and the localization head of our tracker are initialized with weights from ToMP-L, a DINOv2-L variant of ToMP. For an efficient design, the dual model predictors share the same architecture and weights, but two independent lightweight convolutional layers are appended in parallel to the predictors, serving as task-specific heads for semantic weight prediction and perturbation weight prediction, respectively.
We sample 200K subsequences per epoch and train for 30 epochs. Each subsequence consists of two reference frames and one current frame, randomly selected from a 200-frame window within a video sequence. The frames to VGGT are concatenated spatially, which allows better geometric features through multi-frame interaction. Following ToMP (Mayer et al., 2022), PiVOT (Chen et al., 2025a), we set the search area scale factor to $5.0$ and perform data augmentation. The initial learning rate is set to $10^{-4}$ with a StepLR scheduler that decays it by a factor of 0.2 at epochs 10, 15, and 20. AdamW (Loshchilov and Hutter, 2019) is used as the optimizer.
@@ 4.2 Comparisons with the State-of-the-Art Methods
GOT-Edit-378 (Ours) DiNOv2-L VGGT Base+VastTrack 3 53M 64.5 71.1 75.0 80.2* 79.8* 84.8 82.9 75.0 91.0 86.7
@@
ARPTrack-384 (Liang et al., 2025) ViT-ARP-L - Base+VastTrack+K700 7 460M - - - 81.5 80.5 83.4 81.7 74.2 91.1 86.6
@@
GRM-320 (Gao et al., 2023) ViT-MAE-L - Base 3 308M 54.5 66.9 68.9 73.4 70.4 81.2 77.9 71.4 88.9 84.0
SATrack-384 (Ma et al., 2025) SAViT - Base 6 310M 58.4 67.5 - 75.4 73.5 81.4 78.4 72.0 89.0 84.7
@@
GOT-Edit shows a performance gain of about 2–3% across datasets compared with ToMP-378, which is a DINOv2 variant of ToMP (Mayer et al., 2022) and serves as the baseline tracker. Comparing against trackers that employ different semantic backbones, our tracker outperforms all trackers on out-of-distribution targets, except MCITrack-384 (Kang et al., 2025) on in-distribution targets, which uses a different semantic backbone and involves more trainable parameters and frames during training and evaluation. In addition to SUC, NPr, and Pr, we compare trackers using OP50 (Table 3), where all trackers share the same semantic backbone. Our tracker outperforms all others by a clear margin in this metric. We also provide the success AUC curve in Figure 2. On OTB, our method consistently shows the best results. Our tracker outperforms all trackers when $T>0.2$ on AVisT, while outperforming MCITrack when $T<0.7$ on NfS. Additionally, we provide an evaluation on the VOT challenge in Table 2.
@@ 4.3 Ablation Studies
Table 4 presents ablation studies on each GOT-Edit component under image resolution 252, trained using four classical datasets. Row (1) shows the baseline method trained using only semantic features. Row (2) shows that the GOT tracker takes features from the DPT head of VGGT. Even though these features are used to finetune the tracker with GOT data, performance still drops dramatically due to the limited discriminative ability of the geometric information. Row (3) shows the fusion of semantic features from VGGT’s DINO head and geometric features from VGGT’s DPT head, which yields a moderate improvement compared with using geometric features alone. Row (4) shows semantic features extracted from an independent DINO backbone, which perform better than semantic features from the DINO head of VGGT. This is because VGGT fine-tunes its DINO backbone with large-scale 3D data, distorting the original semantic representations of the DINO backbone. Row (5) shows semantic–geometry fusion under the null-space constraint, which improves performance compared with fusion without the constraint. Row (6) shows that whitening and regularization applied to input features before SVD further improve overall performance.
@@ 6 Conclusion
We present GOT-Edit, the first framework to embed geometry-grounded reasoning into generic object tracking via online model editing without explicit 3D inputs. By constraining updates to preserve semantics, GOT-Edit prevents degradation while incorporating geometric cues overlooked by conventional 2D trackers. Through online model editing with null-space constraint, it retains semantic knowledge while adaptively integrating geometric information, achieving robustness under occlusion, clutter, and visual ambiguity. The framework generalizes across datasets, targets, and environments while maintaining stability and robustness. Beyond surpassing state-of-the-art trackers in generalization, the results demonstrate that principled model editing can bridge modality gaps and recover geometry information missed by purely 2D approaches. These advances chart a path toward reliability, safety, and social responsibility in vision systems.
@@
The proposed GOT-Edit framework improves generic object tracking by adaptively integrating semantic and geometric reasoning through online model editing. This capability offers potential societal benefits, including greater reliability of autonomous and robotic systems and improved assistance in challenging visual environments. However, the method may be misused for intrusive surveillance or other applications that compromise privacy and security. Deployment must therefore comply with legal and ethical standards, particularly in contexts involving personal data or sensitive environments. The tracker is trained solely on publicly available datasets, consistent with existing methods and in accordance with established ethical standards. Responsible use requires transparency, rigorous validation, and adherence to established ethical guidelines.
@@
To ensure reproducibility, detailed implementation instructions for GOT-Edit are provided in 4.1. The source code will be publicly available upon acceptance. These measures are intended to facilitate the verification and replication of the results by other researchers.
@@ Appendix A APPENDIX
This document supplements the main paper with details on GOT-Edit, including comparisons with state-of-the-art trackers using NPr, Pr, and SUC plots, ablation on model complexity, and a video appendix for qualitative visualisation on LaSOT and AVisT, available as a zipped file from the paper submission forum.
@@ Appendix C More Experiments
To enhance speed performance, we utilize StreamVGG (Zhuo et al., 2025) to replace VGGT for geometric feature extraction and report the results in Table 8. In this table, ‘GlobalAttn FineTune’ refers to using DoRA (Liu et al., 2024b) to fine-tune the linear layers of the global attention layer in the geometry model, where the global attention layer is the key mechanism for handling cross-frame information. ‘MemCache’ refers to the number of historical K/V caches used for tracking. ‘Frequency’ denotes the frequency for geometric feature extraction. The DoRA rank is set to 16, and only 2.4 M parameters are fine-tuned for the geometry model. The experimental results in the table demonstrate that optimized geometric variants and selective feature application (we set the memory cache to 3 and apply geometric information every 3 frames in the StreamVGGT variant) can significantly increase the speed (e.g., runtime is reduced by approximately 40% when StreamVGGT replaces VGGT, while competitive accuracy is maintained.

### v2

@@ Abstract
Human perception for effective object tracking in a 2D video stream arises from the implicit use of prior 3D knowledge combined with semantic reasoning. In contrast, most generic object tracking (GOT) methods primarily rely on 2D features of the target and its surroundings while neglecting 3D geometric cues, which makes them susceptible to partial occlusion, distractors, and variations in geometry and appearance. To address this limitation, we introduce GOT-Edit, an online cross-modality model editing approach that integrates geometry-aware cues into a generic object tracker from a 2D video stream. Our approach leverages features from a pre-trained Visual Geometry Grounded Transformer to enable geometric cue inference from only a few 2D images. To tackle the challenge of seamlessly combining geometry and semantics, GOT-Edit performs online model editing with null-space constrained updates that incorporate geometric information while preserving semantic discrimination, yielding consistently better performance across diverse scenarios. Extensive experiments on multiple GOT benchmarks demonstrate that GOT-Edit achieves superior robustness and accuracy, particularly under occlusion and clutter, establishing a new paradigm for combining 2D semantics with 3D geometric reasoning for generic object tracking. Project is available [^1] https://chenshihfang.github.io/GOT-EDIT.
@@ 1 Introduction
While geometric information is potentially beneficial for GOT, effectively balancing its contribution with crucial or even dominant semantic information remains a key challenge. As evidenced by our later experiment, a naive fusion strategy improves geometry attributes in tracking but degrades semantic attributes. To address this issue, we propose an online model editing technique that better integrates 3D geometric features (Wang et al., 2025a) with 2D semantic features (Oquab et al., 2023) from only 2D streaming inputs. Our approach is inspired by the null-space model editing of AlphaEdit (Fang et al., 2025), which integrates new knowledge into a trained model while preserving existing knowledge through null-space constraints. However, AlphaEdit performs offline editing, whereas GOT requires online updates to handle dynamically varying targets and backgrounds in seen and unseen scenarios. To bridge this gap, we develop an online editing technique that enables a tracker to adaptively complement 2D semantics with 3D geometric features.
As illustrated in Figure 1, our system begins by extracting both semantic and geometric features from the current and reference frames. These features are then aligned and fused to create an enriched representation, which serves as new knowledge for online tracker adaptation. Built upon the ToMP (Mayer et al., 2022), our approach employs two model predictors: one for the semantic branch and one for the geometric branch. These predictors generate the model weights for the localization head. During tracking, the reference labels, which provide correspondences between the reference frames and serve as few-shot examples of previously predicted and observed information, are dynamically updated to guide the tracker toward the target object. This process guides the model predictors to forecast model weights for the current frame in an online manner. Namely, the semantic model predictor estimates the semantic weights, while the geometry predictor generates complementary weights. A null-space constraint is applied before combining these two sets of weights to preserve the semantic information. Finally, the combined model weights are used by the localization head to localize the target in the current frame.
@@
Figure 1: The GOT-Edit Framework. GOT-Edit facilitates the understanding of 3D geometry to aid generic object tracking from 2D streaming inputs. It predicts semantic and geometric model weights concurrently to incrementally adapt the tracking model. Through online model editing, it ensures geometry-aware, semantic-preserving updates to the tracking model. The solid red box marks the ground-truth target in the input reference frames. The dashed red boxes indicate these same annotations utilized for the online knowledge update within the geometry branch. The green box represents the final predicted tracking result.
@@ 3.2 GOT-Edit
This projector is then utilized in Eq. 10 to compute the geometry-aware perturbation weights.
@@ 4.1 Experimental Setting
Implementation Details. Our method is implemented using PyTorch 2.0.0 and CUDA 11.7. We train the model on eight NVIDIA RTX 4090 GPUs (24 GB each). DeepSpeed (Rasley et al., 2020) is integrated to accelerate training. We also verify that applying activation checkpointing to the tracker further reduces memory consumption, enabling training of the tracker at high resolution (378 $\times$ 378) on four 24 GB GPUs. Inference is performed on a single NVIDIA RTX 4090 GPU and consumes approximately 9 GB of GPU memory during evaluation.
Following PiVOT (Chen et al., 2025a) and LoRAT (Lin et al., 2024), we use the DINOv2 (Oquab et al., 2023) ViT-L backbone for image feature extraction. We initialize the model predictors and localization head with pretrained ToMP-L weights, a ToMP variant with a DINOv2-L backbone. The backbone remains frozen during tracker training. For integrating geometric information, we extract intermediate features from the DPT head of VGGT (Wang et al., 2025a), which is kept frozen during training. We also evaluate an alternative geometry backbone using Depth Anything 3 (Lin et al., 2026), as in Table 1, and StreamVGGT (Zhuo et al., 2026), as in the Appendix.
For an efficient design, the dual model predictors share the same architecture and weights, but two independent lightweight convolutional layers are appended in parallel to the predictors, serving as task-specific heads for semantic weight prediction and perturbation weight prediction, respectively.
We sample 200K subsequences per epoch and train for 25 epochs. Each subsequence consists of two reference frames and one current frame, randomly selected from a 200-frame window within a video sequence. The frames to VGGT are concatenated spatially, which allows better geometric features through multi-frame interaction. Following ToMP (Mayer et al., 2022), PiVOT (Chen et al., 2025a), we set the search area scale factor to $5.0$ and perform data augmentation. The initial learning rate is set to $10^{-4}$ with a StepLR scheduler that decays it by a factor of 0.2 at epochs 10, 15, and 20. AdamW (Loshchilov and Hutter, 2019) is used as the optimizer.
@@ 4.2 Comparisons with the State-of-the-Art Methods
GOT-Edit-378 (Ours) DINOv2-L VGGT Base+VastTrack 3 53M 64.5 71.1 75.0 80.2 79.8 84.8 82.9 75.0 91.0 86.7
@@
GOT-Edit-378 (Ours) DINOv2-L DA3-L Base+VastTrack 3 28M 64.7 70.8 74.8 79.5 79.6 85.2 83.4 75.5 91.0 86.6
GOT-Edit-252 (Ours) DINOv2-L DA3-L Base+VastTrack 3 51M 63.0 70.7 72.6 76.4 75.8 84.5 81.7 73.9 90.1 85.6
@@
ARPTrack-384 (Liang et al., 2025) ViT-ARP-L - Base+VastTrack+K700 7 405M - - - 81.5 80.5 83.4 81.7 74.2 91.1 86.6
@@
GRM-320 (Gao et al., 2023) ViT-MAE-L - Base 3 - 54.5 66.9 68.9 73.4 70.4 81.2 77.9 71.4 88.9 84.0
SATrack-384 (Ma et al., 2025) SAViT - Base 6 - 58.4 67.5 - 75.4 73.5 81.4 78.4 72.0 89.0 84.7
@@
GOT-Edit shows a performance gain of about 2–3% across datasets compared with ToMP-378, which is a DINOv2 variant of ToMP (Mayer et al., 2022) and serves as the baseline tracker. Comparing against trackers that employ different semantic backbones, our tracker outperforms all trackers on out-of-distribution targets, except MCITrack-384 (Kang et al., 2025) on in-distribution targets, which uses a different semantic backbone and involves more trainable parameters and frames during training and evaluation. In addition to SUC, NPr, and Pr, we compare trackers using OP50 (Table 3), where all trackers share the same semantic backbone, and GOT-Edit uses VGGT as the geometry backbone. Our tracker outperforms others by a clear margin on this metric. We also provide the success AUC curve in Figure 2. On OTB, our method consistently shows the best results. Our tracker outperforms all trackers when $T>0.2$ on AVisT, while outperforming MCITrack when $T<0.7$ on NfS. Additionally, we provide an evaluation on the VOT challenge in Table 2.
@@ 4.3 Ablation Studies
Table 4 presents ablation studies on each GOT-Edit component under image resolution 252, trained using four classical datasets. Row (1) shows the baseline method trained using only semantic features. Row (2) shows that the GOT tracker takes features from the DPT head of VGGT. Even though these features are used to finetune the tracker with GOT data, performance still drops dramatically due to the limited discriminative ability of the geometric information. Row (3) shows the fusion of semantic features from VGGT’s DINO head and geometric features from VGGT’s DPT head, which yields a moderate improvement compared with using geometric features alone. Row (4) shows semantic features extracted from an independent DINO backbone, which perform better than semantic features from the DINO head of VGGT. This effect can be largely attributed to the fine-tuning of the DINO backbone of VGGT with large-scale 3D data, which distorts the original semantic representations of the DINO backbone. Row (5) shows semantic–geometry fusion under the null-space constraint, which improves performance compared with fusion without the constraint. Row (6) shows that whitening and regularization applied to input features before SVD further improve overall performance.
@@ 5 Conclusion
We present GOT-Edit, the first framework to integrate 3D geometric cues into generic object tracking via online model editing, while using only 2D streaming inputs during tracking. By constraining updates to preserve semantics, GOT-Edit prevents semantic degradation while incorporating geometric cues that conventional 2D trackers overlook. Through online model editing with null-space constraint, it retains semantic knowledge while adaptively integrating geometric information, achieving robustness under occlusion, clutter, and visual ambiguity. The framework generalizes across datasets, targets, and environments while maintaining stability and robustness. Beyond surpassing state-of-the-art trackers in generalization, the results demonstrate that principled model editing can bridge modality gaps and recover geometry information missed by purely 2D approaches. These advances chart a path toward reliability, safety, and social responsibility in vision systems.
@@
The proposed GOT-Edit framework improves generic object tracking by adaptively integrating semantic and geometric reasoning through online model editing. This capability offers potential societal benefits, including greater reliability of autonomous and robotic systems and improved assistance in challenging visual environments. However, the method may be misused for intrusive surveillance or other applications that compromise privacy and security. Deployment must therefore comply with legal and ethical standards, particularly in contexts involving personal data or sensitive environments. The tracker is trained solely on publicly available datasets, consistent with existing methods and in accordance with established ethical standards. As a pioneering method that enhances tracking by integrating visual geometry, GOT-Edit highlights the emerging, potentially imperfect nature of 3D inference from 2D inputs as a key direction for ongoing technical improvement. Responsible use requires transparency, rigorous validation, and adherence to established ethical guidelines.
@@
To ensure reproducibility, detailed implementation instructions for GOT-Edit are provided in 4.1. The source code will be publicly available. These measures are intended to facilitate the verification and replication of the results by other researchers.
@@ Appendix A APPENDIX
This document supplements the main paper with details on GOT-Edit, including comparisons with state-of-the-art trackers using NPr, Pr, and SUC plots, ablation studies on model complexity, and a video appendix for qualitative visualization, available on the paper submission forum.
@@ Appendix C More Experiments
To enhance speed performance, we utilize StreamVGGT (Zhuo et al., 2026) to replace VGGT for geometric feature extraction and report the results in Table 8. In this table, ‘GlobalAttn FineTune’ refers to using DoRA (Liu et al., 2024b) to fine-tune the linear layers of the global attention layer in the geometry model, where the global attention layer is the key mechanism for handling cross-frame information. ‘MemCache’ refers to the number of historical K/V caches used for tracking. ‘Frequency’ denotes the frequency for geometric feature extraction. The DoRA rank is set to 16, and only 2.4 M parameters are fine-tuned for the geometry model. The experimental results in the table demonstrate that optimized geometric variants and selective feature application (we set the memory cache to 3 and apply geometric information every 3 frames in the StreamVGGT variant) can significantly increase the speed (e.g., runtime is reduced by approximately 40% when StreamVGGT replaces VGGT, while competitive accuracy is maintained.